When an additional charge of 2C is given to a capacitor energy stored in it increasad by 21% .The original charge of the capacitor is 
A.30C
B.40C
​C.10C
D.20C

Grade:12th pass

1 Answers

Arun
5 years ago
Dear student
 
U = q^2 /2c
 
(delU2 – del U1 / del U2) * 100 = [(q + 2 /q)^2 – 1] * 100
 
0.21 = [(q+2 )/q]^2 – 1
 
1.21 = [(q +2)/q]^2
 
q+2 /q = 1.1
 
2/q = .1
 
q = 20